OpenStudy (anonymous):

Find the perimeter and area of a right triangle if the shortest side is 20 m. and the longest side is 29 m. Include correct units with each part of your solution.

OpenStudy (anonymous):

cant anyone help me with this

OpenStudy (anonymous):

Use Pythagorean theorem here..

OpenStudy (allank):

According to the information, we have a right triangle with hypotenuse 29m.

OpenStudy (allank):

Do as @waterineyes says

OpenStudy (anonymous):

\[Third Side = \sqrt{29^2 - 20^2} = \sqrt{441}\] What is the square root of 441 tell me @naomikernizan12

OpenStudy (anonymous):

21

OpenStudy (anonymous):

Yes so third side is 21.. Now can you add all the three sides?? Tell me the result after adding them..

OpenStudy (anonymous):

How?? I am saying to add all the sides: First Side = 21 Second Side = 29 Third Side = 20 What is their sum??

OpenStudy (anonymous):

70

OpenStudy (anonymous):

Yes you are right...

OpenStudy (anonymous):

Basically Perimeter of a Triangle is given by Sum of its all the sides.. Remember this for lifetime..

OpenStudy (anonymous):

thank you and i will :)